OpenStudy (anonymous):

Urgent I have exam 2 morow ln |x-2| = ln | 2x-1| How would u find the inequality and would know if it's a solution or not

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@perl

OpenStudy (perl):

this is not an inequality.

OpenStudy (perl):

want to solve it?

OpenStudy (perl):

solve | x- 2 | = | 2x - 1|

OpenStudy (anonymous):

X-2=2x-1 X= 1

OpenStudy (anonymous):

@Jhannybean can u help me plz

OpenStudy (jhannybean):

seems like perl helped you solve it already?

OpenStudy (anonymous):

No we have to write something about solution such as is it undefine

OpenStudy (anonymous):

I didn't get that part so I know if x = 1 if we substitute that value into back to eq we get -1 and 1 which both are undefine as Ln to 1 is 0 and there are no negative for ln

OpenStudy (anonymous):

Am I right?